air contains 20% by volume of oxygen . the volume of air required for the complete combustion of 1litre of methane under the same conditions is ?
1
Expert's answer
2015-07-28T03:46:34-0400
Reaction of methane combustion is represented below:
CH4 + 3O2 = CO2 + 2H2O
So to burn 1litre of methane we need 3 times higher volume of oxygen, what is 3 liters.
But in the air is just 20% by volume of oxygen, so we will need
V = 3 · 100/20 = 15 litres
Answer: 15 liters of air